Consider the following series 56, 28, 14..







I. Find 17th term.







ii. Find the sum of the series if it continues indefinitely







iii. Find 20th term.

1
Expert's answer
2021-12-05T23:19:10-0500
"\dfrac{28}{56}=\dfrac{1}{2}=\dfrac{14}{28}"

We have an geometric series with "a=56, r=\dfrac{1}{2}."

Since "|r|=\dfrac{1}{2}<1," the geometric series "\displaystyle\sum_{n=0}^{\infin}56(\dfrac{1}{2})^n" converges.


"a_n=ar^{n-1}"

"S=\dfrac{a}{1-r}"

i.

"a_{17}=56(\dfrac{1}{2})^{17-1}=\dfrac{7}{8192}"

ii.


"S=\dfrac{56}{1-\dfrac{1}{2}}=112"

iii.


"a_{20}=56(\dfrac{1}{2})^{20-1}=\dfrac{7}{65536}"